# A Compressive-Expressive Communication Framework for Compositional Representations

### v3

@@ 3.2.2 Pairwise distance maximization (PDM) between messages
$\mathcal{L}^{im}_{\phi^{t+1}}:=\mathcal{L}^{rec}_{\phi^{t+1}}-\beta\frac{1}{N_{batch}^{2}}\sum_{i,j}d(S_{\phi^{t+1}}(x_{i}),S_{\phi^{t+1}}(x_{j})),$ (4)
@@
In § F.4, we show that this regularization term provides a lower bound on entropy maximization objectives [57], and serves as an upper bound for contrastive losses such as NT-Xent [8], thus connecting our formulation to well-established principles in representation learning.
4 Results
@@ 4.1 Datasets
We tested our framework using the Shapes3D [37] dataset, which consists of colored images of 3D geometric shapes, with 6 underlying generating factors $G$: floor hue, wall hue, object hue, shape, scale and orientation. The total amount of attribute-value combinations is $480,000$.
@@
We also evaluated using the compositional split of the MPI3D dataset [17]. This dataset consists of colored images of a robot arm interacting with objects, rendered in controlled 3D scenes with 7 underlying generative factors $G$: object color, object shape, object size, camera height, background color, horizontal arm position, vertical arm position. The dataset comprises $1,036,800$ unique combinations of these factor values.
For both datasets we use the compositional split from Schott et al. [51], which ensures all attribute values appear in training, yet some combinations are reserved for testing set (See §D for a proof that this situation would anyways hold for random train/test splittings with high probability).
@@ 4.2 Evaluation Metrics
Following Chaabouni et al. [6], we assess compositionality using Topographic Similarity (TopSim) [3], a widely used proxy for compositionality. TopSim measures the correlation between pairwise Hamming distances in message space and factor space, capturing the extent to which similar inputs yield similar messages under a structured encoding.
To evaluate communication efficiency, we define the useful length ^$\hat{\ell}_{\epsilon}(x)$ , which estimates the minimal prefix of a message necessary to achieve near-maximal reconstruction quality. Formally:
^$\hat{\ell}_{\epsilon}(x):=\min\left\{i\in\{1,\dots,C\}:\text{MSE}\left(x,R_{\omega}(S_{\phi}(x)_{[i]})\right)\leq\epsilon\right\},$
@@
Finally, to approximate the expressivity of the communication protocol, we evaluate reconstruction quality using the MSE between the generated image and the original input. For all experiments, we report the mean and standard error of 5 seeds.
Additional implementation details are provided in the supplementary material.
@@ 4.3 Progressive decoding reduces the message’s useful length
Table 1: Performance comparison across all experiments on the Shapes3D and MPI3D datasets. We report Topographic Similarity (TopSim $\uparrow$), Useful Length (^$\hat{\ell}_{\epsilon}$downarrow), and Last Symbol MSE ($\downarrow$) for each ablation. The results are grouped by the components of our proposed framework: Interaction, Imitation, and Regularization. Our proposed methods (PD+FiSI+PDM) consistently improve compositionality and message efficiency over the baseline and prior variants.
TopSim $\uparrow$ ^$\hat{\ell}_{2\times 10^{-1}}$downarrow Last symbol MSE $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.259\pm 0.02$ $10.0\pm 0.0$ $0.203\pm 0.03$
Progressive decoding(PD, ours) $0.278\pm 0.03$ $7.8\pm 1.33$ $0.199\pm 0.03$
Imitation Message imitation (baseline) $0.277\pm 0.04$ $9.8\pm 0.4$ $0.189\pm 0.04$
Full state imitation $0.275\pm 0.03$ $10.0\pm 0.0$ $\mathbf{0.154}\pm 0.02$
Final state imitation (FiSI, ours) $0.278\pm 0.05$ $10.0\pm 0.0$ $0.176\pm 0.02$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.265\pm 0.01$ $8.0\pm 1.41$ $0.184\pm 0.03$
PD+FiSI+PDM (ours) $\lambda=1.5$ $\mathbf{0.295\pm 0.03}$ $\mathbf{7.6\pm 1.36}$ $0.182\pm 0.04$
@@
TopSim $\uparrow$ ^$\hat{\ell}_{1.95\times 10^{-2}}$downarrow Last symbol MSE $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.13\pm 0.0$ $9.0\pm 0.28$ $\mathbf{0.014\pm 0.0}$
Progressive decoding(PD, ours) $0.141\pm 0.00$ $\mathbf{6.4\pm 1.31}$ $\mathbf{0.014\pm 0.0}$
Imitation Message imitation (baseline) $0.151\pm 0.01$ $10.0\pm 0.0$ $0.02\pm 0.0$
Full state imitation $0.15\pm 0.03$ $10.0\pm 0.0$ $0.018\pm 0.0$
Final state imitation (FiSI, ours) $0.155\pm 0.01$ $10.0\pm 0.0$ $0.02\pm 0.0$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $\mathbf{0.16\pm 0.0}$ $\mathbf{5.0\pm 1.17}$ $0.02\pm 0.0$
PD+FiSI+PDM (ours) $\lambda=1.5$ $\mathbf{0.16\pm 0.0}$ $7.2\pm 1.53$ $0.02\pm 0.0$
@@
We begin by evaluating the effect of incorporating progressive decoding during the interactive phase of training. As shown in the first two rows of Table 1(b), this modification leads to a 24% reduction in the useful message length, indicating more efficient communication. This gain is achieved without compromising expressivity, evidenced by a slight improvement in reconstruction quality.
To further promote efficient and structured communication, we introduce a geometric penalty term $\lambda$ that penalizes reconstruction error more heavily when longer messages are used. As shown in Figure 2, increasing $\lambda$ consistently improves reconstruction quality, indicating higher expressivity. However, with respect of this parameter, we observe a trade-off between symbolic efficiency and compositionality: while larger values of $\lambda$ encourage more informative messages, they can also lead to less systematic alignment between message space and underlying factors. Intermediate values of $\lambda$ achieve the best balance, maintaining high reconstruction accuracy while preserving TopSim and reducing message length.
@@
Figure 2: Effect of geometric penalty $\lambda$ on emergent communication. We introduce a geometric weighting term $\lambda$ that penalizes reconstruction error more strongly for longer messages. Increasing $\lambda$ improves reconstruction accuracy (b), reflecting enhanced expressivity. Values close to 1 of $\lambda$ strike a favorable balance—achieving efficient, expressive, and structured communication.
@@ 4.4 Final State Imitation increases TopSim
In these experiments, we compare our approach to a baseline model without imitation and to a standard IL variant that imitates the teacher’s message directly—a strategy commonly adopted in prior work for optimizing this phase [42, 43, 34].
Across both the Shapes3D and MPI3D datasets, we find that our final state imitation method consistently enhances the compositionality of the emergent messages. It outperforms the non-IL baseline on both datasets and surpasses message-level imitation on MPI3D. On Shapes3D, it achieves a marginal improvement over message imitation.
More broadly, our results support prior findings that IL promotes greater structure in emergent communication protocols [34, 42]. They also underscore the value of shifting the learning objective from message imitation to image reconstruction, which enforces a tighter generational bottleneck by requiring the student to recover the intended output without access to the original message.
@@ 4.5 Pairwise distance maximization increases compositionality and efficiency
We evaluate a model that combines all previously identified best-performing components—specifically, $\lambda=1.5$, final-state imitation, PD and further introduce an entropy-based regularization term using PDM. We compare this with the KoLeo entropy estimator proposed by Sablayrolles et al. [50].
On the Shapes3D dataset, PDM leads to a marked improvement in TopSim and a lower useful message length, indicating more efficient and structured communication, albeit with a slight increase in reconstruction error. For MPI3D, while this configuration still achieves the highest TopSim, it increases useful length and leads to a marginally higher reconstruction loss compared to using only PD. This highlights a trade-off between expressivity and compressibility, in line with prior theoretical accounts on communication efficiency and compositionality [39, 26, 36].
When comparing PDM to the KoLeo estimator, we find that on Shapes3D, PDM achieves superior TopSim and shorter useful length, with slight reconstruction gains. On MPI3D, KoLeo achieves lower useful length, and both regularizers perform equally in reconstruction and compositionality.
@@ 4.6 Comparison to disentangled representation learning
Following the comparison made by Xu et al. [59], we test our generated messages against standard self-supervised disentangled representation learning frameworks, namely $\beta$-VAE and $\beta$-TCVAE, as well as the baseline VAE+EL[59]. For the continuous models, we extract the first half of the latent vector (corresponding to the predicted means $\mu$) and evaluate it using the DCI disentanglement score [14]. In addition, we train a two-layer MLP to predict ground-truth generative factors, reporting RMSE for continuous factors and classification accuracy for categorical ones.
For the discrete models, we use the predicted messages for each image. We calculate the DCI disentanglement as-is, and use a embedding matrix to transform the discrete messages into the same dimension as the continuous vectors for the MLP. Training is conducted on three subsets of size 1000 of the training set and averaged. Evaluations are conducted on the test sets of the compositional splits for the Shapes3D and MPI3D datasets.
Across both datasets, our method (VAE+CELEBI) demonstrates consistent improvements in disentanglement over baseline methods. On Shapes3D, we observe a substantial increase in DCI scores compared to continuous models and the discrete baseline. However, both discrete models fall behind the continuous baselines in downstream accuracy and RMSE, suggesting a trade-off between representation interpretability and usefulness in downstream tasks.
On MPI3D, our method slightly under-performs when compared to the continuous baselines in categorical accuracy, however greatly outperforming all baselines in continuous regression. We speculate that this discrepancy arises from the lower correlation between pixel-level input statistics and continuous generative factors in MPI3D as opposed to Shapes3D, which may favor communication-based models over purely reconstructive. We provide per-factor evaluation metrics and ablation results in the supplementary material.
Table 2: Disentanglement metrics for Shapes3D and MPI3D datasets.
@@
Shapes3D $\beta$-VAE $0.045\pm 0.004$ $2.42\pm 0.14$ $0.71\pm 0.02$
$\beta$-TCVAE $0.043\pm 0.005$ $\mathbf{2.31\pm 0.12}$ $\mathbf{0.72\pm 0.03}$
VAE + EL $0.108\pm 0.001$ $3.16\pm 0.03$ $0.38\pm 0.03$
VAE + CELEBI $\mathbf{0.110}\pm 0.001$ $3.04\pm 0.08$ $0.42\pm 0.02$
MPI3D $\beta$-VAE $0.031\pm 0.001$ $21.81\pm 0.18$ $\mathbf{0.58\pm 0.01}$
$\beta$-TCVAE $0.031\pm 0.001$ $21.86\pm 0.11$ $0.58\pm 0.01$
VAE + EL $0.108\pm 0.001$ $15.57\pm 0.22$ $0.53\pm 0.00$
VAE + CELEBI $\mathbf{0.136\pm 0.008}$ $\mathbf{14.82\pm 0.06}$ $0.55\pm 0.00$
@@ 5 Related Work
The use of language reference games (LRGs) in conjunction with iterated learning for the purpose of learning self-supervised representations remains underexplored. Xu et al. [59] compare an emergent language autoencoder—comprising a visual backbone and an LSTM-based sender and receiver—with disentanglement frameworks such as the $\beta$-VAE and $\beta$-TCVAE. They find that the representations induced by the emergent language model generalize better in downstream tasks, highlighting the potential of symbolic communication as a compositional bottleneck.
Many previous works have explored using emergent language to induce compositional behavior [23, 13, 1, 12, 5]. Several works utilize a referential LSG in conjunction with IL, using predicted messages as a generational bottleneck [42, 34]. Alternative bottlenecks for emergent language have been proposed, including simplicial embeddings [45, 15], code-books [61] and noisy channels [54].
To replicate the Zipfian distributions observed in natural languages [63, 21], Rita et al. [46] propose the LazImpa framework. Here, an impatient listener produces predictions at every timestep over a variable-length message, and a lazy speaker is explicitly penalized based on message length and prediction accuracy. While both LazImpa and our proposed Progressive Decoding aim to induce incremental informativeness and efficiency by reconstructing the input at every timestep, they differ in key respects. LazImpa operates with variable-length messages and introduces an explicit message length penalty, which we don’t require. In contrast, Progressive Decoding is arguably simpler, as it uses fixed-length messages and, in conjunction with expressivity pressures from our imitation scheme, leads to more topographically structured representations.

### v4

@@ 3.2.2 Pairwise Distance Maximization (PDM)
$\mathcal{L}^{im}_{\phi^{t+1}}:=\mathcal{L}^{rec}_{\phi^{t+1}}+\beta\frac{1}{N_{batch}^{2}}\sum_{i,j}d(S_{\phi^{t+1}}(x_{i}),S_{\phi^{t+1}}(x_{j})),$ (3)
@@
In § F.4, we show that this regularization term provides a lower bound on entropy maximization objectives [entropymax], and serves as an upper bound for contrastive losses such as NT-Xent [chen2020simpleframeworkcontrastivelearning], thus connecting our formulation to well-established principles in representation learning.
4 Experimental setup
@@ 4.1 Datasets
We tested our framework using the Shapes3D [dsprites17] dataset, which consists of colored images of 3D geometric shapes, with 6 underlying generating factors $G$: floor hue, wall hue, object hue, shape, scale, and orientation. The total amount of attribute-value combinations is $480,000$.
@@
We also evaluated using the compositional split of the MPI3D dataset [gondal2019transfer]. This dataset consists of colored images of a robot arm interacting with objects, rendered in controlled 3D scenes with 7 underlying generative factors $G$: object color, object shape, object size, camera height, background color, horizontal arm position, and vertical arm position. The dataset comprises $1,036,800$ unique combinations of these factor values.
For both datasets we use the compositional split from schott2022visual, which ensures all attribute values appear in training, yet some combinations are reserved for the testing set (See §D for a proof that this situation would anyways hold for random train/test splittings with high probability).
@@ 4.2 Evaluation Metrics
Following chaabouni2022emergent, we assess compositionality using Topographic Similarity (TopSim) [Brighton2006UnderstandingLE], a widely used proxy for compositionality. TopSim measures the correlation between pairwise Hamming distances in message space and the generating factor space, capturing the extent to which semantically similar inputs yield similar messages under a structured encoding.
To evaluate communication efficiency, we define the useful length $\hat{\ell}_{\epsilon}(x)$ , which estimates the minimal prefix of a message necessary to achieve near-maximal reconstruction quality. Formally:
$\hat{\ell}_{\epsilon}(x):=\min\left\{i\in\{1,\dots,C\}:\text{MSE}\left(x,R_{\omega}(S_{\phi}(x)_{[i]})\right)\leq\epsilon\right\},$
@@
Finally, to approximate the expressivity of the communication protocol, we evaluate reconstruction quality using the MSE between the generated image and the original input.
We discuss the use of additional language metrics in §I.
5 Results
For all experiments, we report the mean and standard error of 10 random seeds. To assess the robustness of our findings, we also perform permutation tests to evaluate the statistical significance of the observed differences (see §J.1). Additionally, we include qualitative reconstruction experiments in §H, where we visually compare the outputs of all evaluated methods across progressive decoding steps. Further implementation details are provided in §B.
@@ 5.1 Progressive Decoding reduces the message’s useful length
Table 1: Performance comparison across all experiments on the Shapes3D and MPI3D datasets. We report Topographic Similarity (TopSim $\uparrow$), Useful Length ($\hat{\ell}_{\epsilon}$ $\downarrow$), and Last Symbol MSE ($\downarrow$) for each ablation. The results are grouped by the components of our proposed framework: Interaction, Imitation, and Regularization. The synergetic mixture of our proposed methods (PD+FiSI+PDM) consistently improves compositionality and message efficiency over the baseline and prior variants.
TopSim $\uparrow$ $\hat{\ell}_{2\times 10^{-1}}$ $\downarrow$ Last symbol MSE $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.244\pm 0.002$ $10.0\pm 0.0$ $0.212\pm 0.003$
Progressive Decoding(PD, ours) $0.270\pm 0.001$ $7.5\pm 0.215$ $0.238\pm 0.007$
Imitation Message imitation (baseline) $0.257\pm 0.005$ $9.9\pm 0.3$ $0.200\pm 0.007$
Full state imitation $0.256\pm 0.003$ $9.9\pm 0.3$ $0.179\pm 0.003$
Final-State Imitation (FiSI, ours) $0.283\pm 0.003$ $10.0\pm 0.0$ $\mathbf{0.176\pm 0.002}$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.256\pm 0.002$ $8.555\pm 0.167$ $0.179\pm 0.003$
PD+FiSI+PDM (ours) $\lambda=1.5$ $\mathbf{0.292\pm 0.002}$ $\mathbf{7.0\pm 0.155}$ $0.194\pm 0.005$
@@
TopSim $\uparrow$ $\hat{\ell}_{1.95\times 10^{-2}}$ $\downarrow$ Last symbol MSE $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.133\pm 0.001$ $9.3\pm 0.002$ $\mathbf{0.015\pm 0.0}$
Progressive Decoding(PD, ours) $0.137\pm 0.001$ $\mathbf{6.6\pm 0.341}$ $\mathbf{0.015\pm 0.0}$
Imitation Message imitation (baseline) $0.135\pm 0.001$ $9.733\pm 0.029$ $0.016\pm 0.0$
Full state imitation $0.137\pm 0.001$ $9.923\pm 0.020$ $0.018\pm 0.0$
Final-State Imitation (FiSI, ours) $\mathbf{0.156\pm 0.001}$ $9.7\pm 0.046$ $0.02\pm 0.0$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.147\pm 0.002$ $8.9\pm 0.221$ $0.02\pm 0.0$
PD+FiSI+PDM (ours) $\lambda=1.5$ $0.153\pm 0.001$ $9.0\pm 0.167$ $0.02\pm 0.0$
@@
We begin by evaluating the effect of incorporating PD during the interactive phase of training. As shown in the first two rows of Table 1(b), this modification leads to a 25-29% reduction in the useful message length, indicating more efficient communication. These gains are achieved without compromising expressivity on MPI3D and with only a slight compromise on Shapes3D, as indicated by stable reconstruction quality in the former and a minor degradation in the latter.
To further promote efficient and structured communication, we introduce a geometric penalty term $\lambda$ that penalizes reconstruction error more heavily when longer messages are used. As shown in Figure 2, increasing $\lambda$ up to values near 1.5 consistently improves reconstruction quality and useful length, indicating higher expressivity and efficiency. However, when $\lambda$ becomes too large, the reconstruction error incurred at the end of the message dominates the loss function, effectively nullifying PD. We found that setting $\lambda=1.5$ yields favorable results in efficiency and expressivity, while maintaining high compressibility.
@@
Figure 2: Effect of geometric penalty $\lambda$ on emergent communication. We introduce a geometric weighting term $\lambda$ that penalizes reconstruction error more strongly for longer messages. Increasing $\lambda$ initially improves useful length (a). However, once a certain threshold is met, useful length and reconstruction error (b) increase. Values close to 1.5 of $\lambda$ strike a favorable balance—achieving efficient and expressive communication.
@@ 5.2 Final State Imitation increases TopSim
In these experiments, we compare our approach to a baseline model without imitation and to a standard IL variant that imitates the teacher’s message directly—a strategy commonly adopted in prior work for optimizing this phase [ren2020Compositional, ren2024improving, li2019easeofteachinglanguagestructureemergent].
As shown in the imitation section in Table 1(b), in the MPI3D dataset, we find that our final state imitation method consistently enhances the compositionality of the emergent messages. It outperforms the non-IL baseline and surpasses message-level imitation significantly (§ J.1).
More broadly, our results support prior findings that IL promotes greater structure in emergent communication protocols [li2019easeofteachinglanguagestructureemergent, ren2020Compositional]. They also underscore the value of shifting the learning objective from message imitation to image reconstruction, which enforces a tighter generational bottleneck by requiring the student to recover the intended output without access to the original message.
@@ 5.3 Pairwise Distance Maximization increases compositionality and efficiency
We evaluate a model that combines all previously identified best-performing components—specifically, $\lambda=1.5$, final-state imitation, PD and further introduce an entropy-based regularization term using PDM. We compare this with the KoLeo entropy estimator proposed by sablayrolles2018spreading.
When examining the effect of incorporating PDM, Table 1(b) shows that PDM outperforms both FiSi and PD in fostering structure in the message. On the Shapes3D dataset, PDM leads to a marked improvement in TopSim and a lower useful message length, indicating more efficient and structured communication, albeit with a slight increase in reconstruction error. For MPI3D, this configuration doesn’t achieve the highest TopSim, albeit close to the best, and it increases useful length and leads to a marginally higher reconstruction loss compared to using only PD. This highlights a trade-off between expressivity and compressibility, in line with prior theoretical accounts on communication efficiency and compositionality [ambiguityandefficiency, iteratedlearningkirby, Luna2020InternalAE].
When comparing PDM to the KoLeo estimator, Regularization section in Table 1(b), we find that on Shapes3D, PDM achieves superior TopSim and shorter useful length, with slight reconstruction gains. On MPI3D, KoLeo leads to a lower useful length but exhibits reduced compositionality, with both regularizers showing similar reconstruction performance.
@@ 5.4 Comparison to disentangled representation learning
Following the comparison made by elvae, we test our generated messages against standard self-supervised disentangled representation learning frameworks, namely $\beta$-VAE and $\beta$-TCVAE, as well as the baseline VAE+EL [elvae]. For the continuous models, we extract the first half of the latent vector (corresponding to the predicted means $\mu$) and evaluate it using the DCI disentanglement score [DCI2018]. In addition, we train a two-layer MLP to predict ground-truth generative factors, reporting RMSE for continuous factors and classification accuracy for categorical ones.
For the discrete models, we use the predicted messages for each image. We calculate the DCI disentanglement as-is, and use an embedding matrix to transform the discrete messages into the same dimension as the continuous vectors for the MLP. Training is conducted on three subsets of size 1000 of the training set and averaged. Evaluations are conducted on the test sets of the compositional splits for the Shapes3D and MPI3D datasets.
Across both datasets, our method (VAE+CELEBI) demonstrates consistent improvements in disentanglement over baseline methods. On Shapes3D, we observe a substantial increase in DCI scores compared to continuous models and the discrete baseline (see statistical significance tests in § J.2). However, both discrete models fall behind the continuous baselines in downstream accuracy and RMSE, suggesting a trade-off between representation interpretability and usefulness in downstream tasks, which is consistent with previous work indicating that discrete representations may be less accessible to simple classifiers (as shown in [xu2022compositional]), and that inductive biases for compositionality do not imply disentanglement [montero2021the].
On MPI3D, our method slightly under-performs when compared to the continuous baselines in categorical accuracy, however greatly outperforming all baselines in continuous regression. We speculate that this discrepancy arises from the lower correlation between pixel-level input statistics and continuous generative factors in MPI3D as opposed to Shapes3D, which may favor communication-based models over purely reconstructive.
Table 2: Disentanglement metrics for the Shapes3D and MPI3D datasets. We report the DCI score to evaluate disentanglement quality, and the RMSE and classification accuracy of a linear probe trained on the latent representations (both message and vector embeddings) to assess their suitability for downstream tasks.
@@
Shapes3D $\beta$-VAE $0.045\pm 0.004$ $2.460\pm 0.144$ $0.702\pm 0.025$
$\beta$-TCVAE $0.043\pm 0.005$ $\mathbf{2.378\pm 0.123}$ $\mathbf{0.721\pm 0.028}$
VAE + EL $0.108\pm 0.000$ $3.168\pm 0.027$ $0.343\pm 0.019$
VAE + CELEBI $\mathbf{0.112\pm 0.001}$ $2.932\pm 0.045$ $0.453\pm 0.024$
MPI3D $\beta$-VAE $0.031\pm 0.001$ $21.819\pm 0.111$ $0.581\pm 0.006$
$\beta$-TCVAE $0.031\pm 0.001$ $21.854\pm 0.188$ $\mathbf{0.582\pm 0.005}$
VAE + EL $0.114\pm 0.002$ $15.971\pm 0.301$ $0.529\pm 0.006$
VAE + CELEBI $\mathbf{0.137\pm 0.002}$ $\mathbf{14.82\pm 0.012}$ $0.542\pm 0.002$
@@ 6 Related Work
The use of language reference games (LRGs) in conjunction with iterated learning for the purpose of learning self-supervised representations remains underexplored. elvae compare an emergent language autoencoder—comprising a visual backbone and an LSTM-based sender and receiver—with disentanglement frameworks such as the $\beta$-VAE and $\beta$-TCVAE. They find that the representations induced by the emergent language model generalize better in downstream tasks, highlighting the potential of symbolic communication as a compositional bottleneck.
Many previous works have explored using emergent language to induce compositional behavior [kharitonov2020entropy, Dess2021InterpretableAC, auersperger2022defendingcompositionalityemergentlanguages, Denamganai2020OnS, chaabouni2020compositionality]. Several works utilize a referential LSG in conjunction with IL, using predicted messages as a generational bottleneck [ren2020Compositional, li2019easeofteachinglanguagestructureemergent]. Alternative bottlenecks for emergent language have been proposed, including simplicial embeddings [ren2024improving, Elberg_2024], code-books [zheng2024iterated] and noisy channels [ueda2024lewisssignalinggamebetavae]. Our message tokens can also be considered as roughly comparable to the slots in slot-attention architectures such as [eslami2016attend, greff2019multi, Lin2020SPACE, jiangscalor, lin2020improving, Engelcke2020GENESIS, locatello2020object, brady2025interaction, akan2025slotguided], however note that among other differences, the slot-attention typically additive decoding layer [burgess2019monet, lachapelle2023additive] does not compare directly to our proposed Progressive Decoding.
To replicate the Zipfian distributions observed in natural languages [Zipf49, Kanwal2017-hs], rita-etal-2020-lazimpa propose the LazImpa framework. The LazImpa Impatient listener loss is similar to our proposed Progressive Decoding, in that they both aim to induce incremental informativeness and efficiency by reconstructing the input at every timestep. However LazImpa is optimized for a referential game task (and not a reconstruction task as ours) and uses a linear length penalization in conjunction with a cross-entropy loss. Our proposal instead uses the actual reconstruction loss and the length penalization is exponential and based on a tunable parameter $\lambda$.